# Leveraging Multimodal LLM Descriptions of Activity for Explainable Semi-Supervised Video Anomaly Detection

### v1

@@ 1 Introduction
Method Year Focus Features Task Objective
MLLM Interaction Explain- Semi Scene
Usage Anomalies ability Supervised Specific
VadCLIP  [47] 2023 Using CLIP for multi-scene weakly supervised VAD
Vera  [50] 2025 Optimizing prompts for MLLM-based weakly supervised VAD
Holmes-VAU  [53] 2025 MLLM finetuning for anomaly understanding across multi-scene
AnomalyRuler  [46] 2024 Scene-specific rule generation and frame-level descriptions via MLLMs
VLAVAD  [16] 2024 Using frame level MLLM descriptions for semi-supervised VAD
EVAL  [36] 2023 Using object-level features for explainable VAD
T-EVAL  [37] 2024 Object-level, tracklet-based explainable VAD
Scene-Graph  [28] 2025 Exemplar and scene graph for interaction-based VAD
Ours 2026 Object-level MLLM descriptions for interaction-based VAD
@@
Video anomaly detection (VAD) has emerged as an active research field due to its significant applications in security and public safety. The rapid growth of surveillance video data makes it impractical for humans to monitor thoroughly. Consequently, VAD algorithms play a vital role in identifying unusual events in video footage, enabling human operators to focus their attention on reviewing these flagged incidents. In this paper, we focus on the formulation in which nominal videos (also called training videos) containing only normal activities in a particular scene are provided for learning a model. The goal is to identify when and where anomalous activities happen within test videos of the same scene. Along with other papers [3, 44], we refer to this version of the problem as semi-supervised VAD because of the availability of normal, but not anomalous, training video. In Table 1, we summarize the properties of some recent papers in this area to highlight the differences between our work and other recent approaches.
@@
Existing semi-supervised video anomaly detection methods suffer from several weaknesses. Recent studies struggle to provide explainability for detected anomalies. While some existing work [36, 37] explored more explainable approaches by interpreting the input features, they do not offer direct, textual explanations. Furthermore, the ComplexVAD dataset [28] introduced interaction-based anomalies and demonstrated that detecting these anomalies is even more challenging than identifying the simpler anomalies presented in earlier datasets. Hence, we propose a new method to address the drawbacks present in existing works, including the lack of focus on complex anomalies and explainability. By utilizing a Large Language Model-based detection approach, we introduce an effective video anomaly detection framework that excels at identifying complex anomalies and features built-in explainability.
@@
Large Language Models, especially in their multimodal forms (MLLMs), have demonstrated strong capabilities in interpreting and accurately describing complex situations from both text and images. GPT-4 with vision [1], for example, can analyze visual inputs and produce detailed, contextually relevant descriptions. Existing work has explored the use of MLLMs for video anomaly detection. Most of these methods are designed for a multi-scene, weakly supervised setting. While such approaches are effective for detecting a common set of anomalies across multiple scenes, they fall short in identifying scene-specific anomalies, which are central to the semi-supervised, single-scene VAD task. To address this limitation, we propose a novel MLLM-based framework that: (1) considers object interactions in addition to individual object activities, (2) leverages an MLLM to generate high-level descriptions used in an exemplar-based model that can be easily extended without retraining deep networks for each new scene, and (3) enables both spatial and temporal localization of anomalies.
@@
In our method, we adopt an object-level feature extraction strategy similar to [28], using an object detector and tracker. We then employ an MLLM agent to describe object activities and interactions by querying it with pairs of frames, separated in time by a fixed interval, to capture temporal information. A key novelty is using these MLLM-generated textual descriptions directly as features for modeling normal activity. By removing redundant descriptions (Section 3.2), we construct a final set of textual exemplars representing nominal videos. This exemplar set is then used to detect anomalies: descriptions from anomalous interactions are expected to deviate from nominal ones. In summary, our contributions are:
@@ 2 Related Work
Many different formulations of the video anomaly detection (VAD) problem have been studied in recent years: semi-supervised [13, 15, 19, 42], weakly supervised [38], unsupervised [51], training-free [52], single-scene [32] and multi-scene [22]. Each of these formulations has applicability in different scenarios. This paper focuses on the single-scene, semi-supervised version of video anomaly detection because it corresponds to the most common use case that we have observed in the real-world; namely, that of a surveillance camera at a particular location in the world in which video of normal activity is abundant but video containing anomalies is rare and expensive to collect. Furthermore, what is anomalous in one scene may be normal in another. This along with the fact that location-dependent anomalies (e.g. a pedestrian may be anomalous in some locations but not others) are common in the single-scene version of VAD but do not occur in the multi-scene version of VAD implies that the multi-scene version of VAD is not a generalization of single-scene VAD [32]. Solutions for one version generally do not work well on the other version.
Early work on the single-scene, semi-supervised version of VAD includes many papers based on frame reconstruction [13, 15] or frame prediction [19, 42]. One big disadvantage of such approaches is that they require training a deep network on video from each new scene. To avoid this disadvantage, others [8, 7, 31, 30, 36, 37, 28] used an exemplar-based model that extracts a set of representative features from normal video but does not require any deep network training for each scene. We also use this idea in our work.
@@ 2.1 MLLMs in Video Anomaly Detection
Weakly-Supervised and Training-Free VAD. Many approaches to VAD that use an MLLM try to use its built-in knowledge of what is normal instead of learning normality from normal training video [52, 50, 23]. A disadvantage of such approaches is that they cannot handle scenarios such as a boxing gym in which activity that is typically rare and anomalous (fighting) is actually very common. Another important difference between our work and most other MLLM-based approaches to VAD is that others have focused on the weakly supervised, multi-scene version of VAD [17, 45, 52, 50, 23].
Semi-Supervised VAD. Compared to the weakly-supervised setting, the capabilities of MLLMs are understudied in the semi-supervised setting. Wu et al. [46] addresses semi-supervised, single-scene VAD, but with a fundamentally different methodology. Their approach uses an MLLM to first generate a set of textual rules from normal video and then applies these rules during inference. In contrast, our method directly compares MLLM-generated text descriptions between normal and test videos without an intermediate rule-generation step. Furthermore, their model processes full video frames, whereas we employ an object-centric approach that analyzes regions around objects and their pairs. This key design choice allows our method to model object interactions and perform both spatial and temporal localization, while their approach is limited to temporal localization only.
Li and Jiang [16] also focuses on semi-supervised single scene anomaly detection and proposes an MLLM-based method. Like our method, they first detect and track objects and then query an MLLM to yield high-level descriptions of objects and their activity. However, they do not model object interactions, and do not use an exemplar-based model.
In real-world scenarios, many anomalies are characterized by an unusual interaction between two objects. Therefore, one of the important aspects of our method is its modeling of object interactions in the scene to identify anomalous interactions. There have been a few other methods that also modeled interactions [5, 9, 39, 40, 28]. While there are many differences among these methods, one of the main differences compared to our current work is that none of them take advantage of MLLMs. In our approach, the textual descriptions of object activity and interactions provided by an MLLM yield powerful cues for a high-level understanding of a scene.
@@ 3.2 Generating Textual Descriptions
An additional “system” prompt is also used with the MLLM: “You will be provided with two frames from a video and asked to describe what objects that are indicated by bounding boxes in the video are doing. Your task is to answer the query in a simple sentence. If there is any interaction between the indicated objects, a description of the interaction should be given.”
@@ 3.3 Model Building
We follow a similar exemplar selection strategy as introduced in recent works [30, 36, 37, 28]. For all frames in a video, we collect all textual descriptions generated from object pairs into one set and all textual descriptions generated from single objects into another set. Then for each of the sets we run an exemplar selection algorithm which selects a subset of the elements of the set such that no two members of the subset are near each other according to a distance function. The intuition behind exemplar selection is to simply remove redundant (or nearly redundant) elements from the set leaving behind a compact, representative subset of exemplars. Various distance functions between sentences can be used to compare textual descriptions. In most of our experiments we use Sentence-BERT [33], but we also test using BLEU [29] and METEOR [2].
@@ 3.5 Combination with Previous Methods
In our experiments, we show results when combining our method separately with two other methods; namely, the scene-graph method of Mumcu et al. [28] and the tracklet method of Singh et al [37].
Our method of modeling interactions was inspired by the scene-graph method of Mumcu et al. [28]. Their method builds a scene graph for each frame where each object is a node and objects that are nearby and, therefore possibly interacting, are connected by an edge. The attributes stored at each node are the object’s location, size, object class, trajectory and skeletal pose. Then an exemplar-based model is created for every connected pair of nodes across all scene graphs in all nominal frames. A separate exemplar-based model is also created for all singleton nodes across all scene graphs in all nominal frames. To combine our method with theirs, we can simply add the MLLM-generated textual descriptions to each connected pair of nodes and to each singleton node. The distance function used to compare nodes takes the maximum over all attribute distances including the distance between textual descriptions.
In the case of the tracklet method of Singh et al. [37], objects in each frame are also detected and tracked. Similar to Mumcu et al., for each object, a set of attributes are stored including the location, size, object class and trajectory. An exemplar-based model is then learned from all of the objects found in the nominal training videos using a distance function for comparing objects that takes the maximum over each attribute distance. To combine our method with theirs, we again only need to add our MLLM-generated textual descriptions as another attribute associated with each object. In this case, object interactions are not modeled. We use this combination to test on datasets that do not contain anomalous object interactions.
@@ 4 Experiments
Datasets: We evaluate MLLM-EVAD on three video anomaly detection benchmark datasets which are all publicly available under a CC-BY-SA-4.0 license. ComplexVAD [28] is a large-scale dataset with a total of over 3.6 million frames, with 2,069,941 frames for training and 1,611,497 for testing. It features interaction-based anomalies and serves as our main focus, since our method leverages descriptions of object activity and interactions for video anomaly detection. The scene features a crosswalk, pedestrian sidewalks, and a two-lane street, which are heavily used. It demonstrates interaction anomalies such as a person leaving an object on the ground, a person sitting on a car, or a dog walking without a person or leash, where the individual objects are normal, but their interactions are anomalous.
@@
In addition, we evaluate performance on the Avenue and Street Scene datasets. Avenue [21] contains short surveillance videos of a campus walkway, with anomalies such as running, throwing objects, or walking in the wrong direction. Street Scene [30] includes longer videos of a city street environment, where anomalies involve vehicles or pedestrians behaving unusually, including cars driving on the wrong side of the road, pedestrians crossing at non-designated areas, or bikers riding on the sidewalk. Avenue contains 30,652 frames, split nearly evenly between 15,328 training frames and 15,324 testing frames. Street Scene is a larger dataset with a total of 203,257 frames, of which 56,847 are used for training and 146,410 for testing.
@@
Implementation Details: We use Detectron2 [48] and ByteTrack [54] as the object detector and object tracker in our method. Sentence-BERT [33] (using the pretrained weights, ”all-MiniLM-L6-v2”) is used as the text embedding model to encode activity descriptions. Initially, GPT-4o [14] was used as the multi-modal MLLM agent. However, with the recent introduction of Gemma 3 [41], we re-conducted the experiments on ComplexVAD using Gemma 3. Hence, the results on ComplexVAD are reported with both Gemma 3 and GPT-4o, while the results on the other datasets are reported with GPT-4o. We choose a threshold $th=0.65$ for exemplar selection that resulted in a modest number of total exemplars selected. From past work that used exemplar-based models, this threshold mainly effects model size and has a small effect on test accuracy [36, 37]. We provide additional implementation details in the Appendix.
@@
Evaluation Criteria: We use the Region-Based Detection Criterion (RBDC) and the Track-Based Detection Criterion (TBDC) as proposed in [30] as our primary evaluation criteria and report the area under the curve (AUC) for false positive rates per frame from 0 to 1. We also report frame-level AUC [24] scores. As highlighted in previous works [30] frame-level AUC only evaluates temporal accuracy and disregards spatial localization of anomalies. RBDC and TBDC measure a method’s capacity to accurately identify anomalous spatio-temporal regions within a given video sequence.

### v2

@@ 1 Introduction
Video anomaly detection (VAD) has emerged as an active research field due to its significant applications in security and public safety. The rapid growth of surveillance video data makes it impractical for humans to monitor thoroughly. Consequently, VAD algorithms play a vital role in identifying unusual events in video footage, enabling human operators to focus their attention on reviewing these flagged incidents. In this paper, we focus on the formulation in which nominal videos (also called training videos) containing only normal activities in a particular scene are provided for learning a model. The goal is to identify when and where anomalous activities happen within test videos of the same scene. Along with other papers (Baradaran and Bergevin, 2024; Wu et al., 2024a), we refer to this version of the problem as semi-supervised VAD because of the availability of normal, but not anomalous, training video. In Table 1, we summarize the properties of some recent papers in this area to highlight the differences between our work and other recent approaches.
@@
Existing semi-supervised video anomaly detection methods suffer from several weaknesses. Recent studies struggle to provide explainability for detected anomalies. While some existing work (Singh et al., 2023; 2024) explored more explainable approaches by interpreting the input features, they do not offer direct, textual explanations. Furthermore, the ComplexVAD dataset (Mumcu et al., 2025b) introduced interaction-based anomalies and demonstrated that detecting these anomalies is even more challenging than identifying the simpler anomalies presented in earlier datasets. Hence, we propose a new method to address the drawbacks present in existing works, including the lack of focus on complex anomalies and explainability. By utilizing a Large Language Model-based detection approach, we introduce an effective video anomaly detection framework that excels at identifying complex anomalies and features built-in explainability.
@@
Large Language Models, especially in their multimodal forms (MLLMs), have demonstrated strong capabilities in interpreting and accurately describing complex situations from both text and images. GPT-4 with vision (Achiam et al., 2023), for example, can analyze visual inputs and produce detailed, contextually relevant descriptions. Existing work has explored the use of MLLMs for video anomaly detection. Most of these methods are designed for a multi-scene, weakly supervised setting. While such approaches are effective for detecting a common set of anomalies across multiple scenes, they fall short in identifying scene-specific anomalies, which are central to the semi-supervised, single-scene VAD task. To address this limitation, we propose a novel MLLM-based framework that: (1) considers object interactions in addition to individual object activities, (2) leverages an MLLM to generate high-level descriptions used in an exemplar-based model that can be easily extended without retraining deep networks for each new scene, and (3) enables both spatial and temporal localization of anomalies.
@@
In our method, we adopt an object-level feature extraction strategy similar to (Mumcu et al., 2025b), using an object detector and tracker. We then employ an MLLM agent to describe object activities and interactions by querying it with pairs of frames, separated in time by a fixed interval, to capture temporal information. A key novelty is using these MLLM-generated textual descriptions directly as features for modeling normal activity. By removing redundant descriptions (Section 3), we construct a final set of textual exemplars representing nominal videos. This exemplar set is then used to detect anomalies: descriptions from anomalous interactions are expected to deviate from nominal ones. In summary, our contributions are:
@@
Method Focus Features Task Objective
MLLM Interaction Explain- Semi Scene
Usage Anomalies ability Supervised Specific
VadCLIP  Wu et al. (2024d) Using CLIP for multi-scene weakly supervised VAD
Vera  Ye et al. (2025) Optimizing prompts for MLLM-based weakly supervised VAD
Holmes-VAU  Zhang et al. (2025) MLLM finetuning for anomaly understanding across multi-scene
AnomalyRuler  Wu et al. (2024b) Scene-specific rule generation and frame-level descriptions via MLLMs
VLAVAD  Li and Jiang (2024) Using frame level MLLM descriptions for semi-supervised VAD
EVAL  Singh et al. (2023) Using object-level features for explainable VAD
T-EVAL  Singh et al. (2024) Object-level, tracklet-based explainable VAD
Scene-Graph  (Mumcu et al., 2025b) Exemplar and scene graph for interaction-based VAD
Ours Object-level MLLM descriptions for interaction-based VAD
@@ 2 Related Work
Many different formulations of the video anomaly detection (VAD) problem have been studied in recent years: semi-supervised (Hasan et al., 2016; Ionescu et al., 2019; Liu et al., 2018; Wang et al., 2021), weakly supervised (Sultani et al., 2018), unsupervised (Zaheer et al., 2022), training-free (Zanella et al., 2024), single-scene (Ramachandra et al., 2020a) and multi-scene (Luo et al., 2017). Each of these formulations has applicability in different scenarios. This paper focuses on the single-scene, semi-supervised version of video anomaly detection because it corresponds to the most common use case that we have observed in the real-world; namely, that of a surveillance camera at a particular location in the world in which video of normal activity is abundant but video containing anomalies is rare and expensive to collect. An important attribute of the single-scene VAD problem formulation is that anomalies are scene dependent, i.e., what is anomalous in one scene may be normal in another. This is in contrast to the multi-scene problem formulation in which anomaly types are common across different scenes. Finally, the single-scene VAD formulation typically contains anomalies that are location-dependent (e.g. a pedestrian may be anomalous in some locations but not others). Such location-dependent anomalies do not occur in multi-scene VAD because they require scene-specific models. These differences imply that the multi-scene version of VAD is not a generalization of single-scene VAD (Ramachandra et al., 2020a). Solutions for one version generally do not work well on the other version.
Early work on the single-scene, semi-supervised version of VAD includes many papers based on frame reconstruction (Hasan et al., 2016; Ionescu et al., 2019) or frame prediction (Liu et al., 2018; Wang et al., 2021). One big disadvantage of such approaches is that they require training a deep network on video from each new scene. To avoid this disadvantage, others (Doshi and Yilmaz, 2020b; a; Ramachandra et al., 2020b; Ramachandra and Jones, 2020; Singh et al., 2023; 2024; Mumcu et al., 2025b) used an exemplar-based model that extracts a set of representative features from normal video but does not require any deep network training for each scene. We also use this idea in our work.
@@ 2.1 MLLMs in Video Anomaly Detection
Weakly-Supervised and Training-Free VAD. Many approaches to VAD that use an MLLM try to use its built-in knowledge of what is normal instead of learning normality from normal training video (Zanella et al., 2024; Ye et al., 2025; Lv and Sun, 2024). A disadvantage of such approaches is that they cannot handle scenarios such as a boxing gym in which activity that is typically rare and anomalous (fighting) is actually very common. Another important difference between our work and most other MLLM-based approaches to VAD is that others have focused on the weakly supervised, multi-scene version of VAD (Li et al., 2025; Wu et al., 2024c; Zanella et al., 2024; Ye et al., 2025; Lv and Sun, 2024).
Semi-Supervised VAD. Compared to the weakly-supervised setting, the capabilities of MLLMs are understudied in the semi-supervised setting.  Wu et al. (2024b) addresses semi-supervised, single-scene VAD, but with a fundamentally different methodology. Their approach uses an MLLM to first generate a set of textual rules from normal video and then applies these rules during inference. In contrast, our method directly compares MLLM-generated text descriptions between normal and test videos without an intermediate rule-generation step. Furthermore, their model processes full video frames, whereas we employ an object-centric approach that analyzes regions around objects and their pairs. This key design choice allows our method to model object interactions and perform both spatial and temporal localization, while their approach is limited to temporal localization only.
Li and Jiang (2024) also focuses on semi-supervised single scene anomaly detection and proposes an MLLM-based method. Like our method, they first detect and track objects and then query an MLLM to yield high-level descriptions of objects and their activity. However, they do not model object interactions, and do not use an exemplar-based model.
In real-world scenarios, many anomalies are characterized by an unusual interaction between two objects. Therefore, one of the important aspects of our method is its modeling of object interactions in the scene to identify anomalous interactions. There have been a few other methods that also modeled interactions (Chen et al., 2018; Doshi and Yilmaz, 2023; Sun et al., 2020; Szymanowicz et al., 2021; Mumcu et al., 2025b). While there are many differences among these methods, one of the main differences compared to our current work is that none of them take advantage of MLLMs. In our approach, the textual descriptions of object activity and interactions provided by an MLLM yield powerful cues for a high-level understanding of a scene. Language provides a natural abstraction for compositional structure and relational semantics (e.g., interactions, roles, and intent), which are difficult to capture reliably with low-level features alone. In the one-class setting, this abstraction enables a compact and human-aligned representation of normal behavior, where deviations correspond to semantically meaningful changes rather than pixel-level noise.
@@ 3.1 Object Detection and Tracking
We use a temporal gap of approximately one second between the paired frames as a practical trade-off between capturing meaningful temporal changes and maintaining object identity consistency in a static-camera, single-scene setting. Shorter temporal gaps often result in nearly identical visual content and textual descriptions with limited motion cues, while significantly larger gaps increase the risk of object drift, missed interactions, or changes in object visibility. Importantly, the proposed framework is not tightly coupled to a specific temporal gap; the paired-frame formulation serves as a lightweight mechanism for incorporating local temporal context, and similar behavior can be captured with other reasonable temporal offsets depending on the scene dynamics.
@@ 3.2 Generating Textual Descriptions
An additional “system” prompt is also used with the MLLM: “You will be provided with two frames from a video and asked to describe what objects that are indicated by bounding boxes in the video are doing. Your task is to answer the query in a simple sentence. If there is any interaction between the indicated objects, a description of the interaction should be given”
@@
We note that our use of image-based multimodal language models is a deliberate design choice. During development, we explored video-language models that process short clips or frame sequences; however, we observed that such models tend to produce coarse, scene-level descriptions that are less sensitive to fine-grained object interactions and localized activity changes. In contrast, querying an image-based MLLM with paired object-centric crops from temporally separated frames provides sufficient temporal context while preserving spatial grounding and object identity. This design aligns naturally with our goal of scene-specific modeling of normal behavior through localized object interactions in a semi-supervised setting, rather than general video understanding.
@@ 3.3 Model Building
We follow a similar exemplar selection strategy as introduced in recent works (Ramachandra and Jones, 2020; Singh et al., 2023; 2024; Mumcu et al., 2025b). For all frames in a video, we collect all textual descriptions generated from object pairs into one set and all textual descriptions generated from single objects into another set. Then for each of the sets we run an exemplar selection algorithm which selects a subset of the elements of the set such that no two members of the subset are near each other according to a distance function. The intuition behind exemplar selection is to simply remove redundant (or nearly redundant) elements from the set leaving behind a compact, representative subset of exemplars. Various distance functions between sentences can be used to compare textual descriptions. In most of our experiments we use Sentence-BERT (Reimers and Gurevych, 2019), but we also test using BLEU (Papineni et al., 2002) and METEOR (Banerjee and Lavie, 2005).
@@ 3.5 Combination with Previous Methods
In our experiments, we show results when combining our method separately with two other methods; namely, the scene-graph method of  Mumcu et al. (2025b) and the tracklet method of  Singh et al. (2024).
Our method of modeling interactions was inspired by the scene-graph method of  Mumcu et al. (2025b). Their method builds a scene graph for each frame where each object is a node and objects that are nearby and, therefore possibly interacting, are connected by an edge. The attributes stored at each node are the object’s location, size, object class, trajectory and skeletal pose. Then an exemplar-based model is created for every connected pair of nodes across all scene graphs in all nominal frames. A separate exemplar-based model is also created for all singleton nodes across all scene graphs in all nominal frames. To combine our method with theirs, we can simply add the MLLM-generated textual descriptions to each connected pair of nodes and to each singleton node. The distance function used to compare nodes takes the maximum over all attribute distances including the distance between textual descriptions.
In the case of the tracklet method of  Singh et al. (2024), objects in each frame are also detected and tracked. Similar to  Mumcu et al. (2025b), for each object, a set of attributes are stored including the location, size, object class and trajectory. An exemplar-based model is then learned from all of the objects found in the nominal training videos using a distance function for comparing objects that takes the maximum over each attribute distance. To combine our method with theirs, we again only need to add our MLLM-generated textual descriptions as another attribute associated with each object. In this case, object interactions are not modeled. We use this combination to test on datasets that do not contain anomalous object interactions.
@@ 4 Experiments
Datasets: We evaluate MLLM-EVAD on three video anomaly detection benchmark datasets which are all publicly available under a CC-BY-SA-4.0 license. ComplexVAD (Mumcu et al., 2025b) is a large-scale dataset with a total of over 3.6 million frames, with 2,069,941 frames for training and 1,611,497 for testing. It features interaction-based anomalies and serves as our main focus, since our method leverages descriptions of object activity and interactions for video anomaly detection. The scene features a crosswalk, pedestrian sidewalks, and a two-lane street, which are heavily used. It demonstrates interaction anomalies such as a person leaving an object on the ground, a person sitting on a car, or a dog walking without a person or leash, where the individual objects are normal, but their interactions are anomalous.
@@
In addition, we evaluate performance on the Avenue and Street Scene datasets. Avenue (Lu et al., 2013) contains short surveillance videos of a campus walkway, with anomalies such as running, throwing objects, or walking in the wrong direction. Street Scene (Ramachandra and Jones, 2020) includes longer videos of a city street environment, where anomalies involve vehicles or pedestrians behaving unusually, including cars driving on the wrong side of the road, pedestrians crossing at non-designated areas, or bikers riding on the sidewalk. Avenue contains 30,652 frames, split nearly evenly between 15,328 training frames and 15,324 testing frames. Street Scene is a larger dataset with a total of 203,257 frames, of which 56,847 are used for training and 146,410 for testing.
@@
Implementation Details: We use Detectron2 (Wu et al., 2019) and ByteTrack (Zhang et al., 2022) as the object detector and object tracker in our method. Sentence-BERT (Reimers and Gurevych, 2019) (using the pretrained weights of all-MiniLM-L6-v2 variant) is used as the text embedding model to encode activity descriptions. Initially, GPT-4o (Hurst et al., 2024) was used as the multi-modal MLLM agent. However, with the recent introduction of Gemma 3 (Team et al., 2025), we re-conducted the experiments on ComplexVAD using Gemma 3. Hence, the results on ComplexVAD are reported with both Gemma 3 and GPT-4o, while the results on the other datasets are reported with GPT-4o. We choose a threshold $th=0.65$ for exemplar selection, which results in a modest number of total exemplars while maintaining strong detection performance. As observed in prior exemplar-based VAD methods, the choice of $th$ primarily affects model size rather than test accuracy, with performance remaining stable across a reasonable range of threshold values (Singh et al., 2023; 2024). In our framework, exemplar selection serves to remove near-duplicate descriptions and construct a compact representation of normal behavior, and anomaly detection relies on relative distances to this representation rather than precise threshold tuning. We provide additional implementation details in the Appendix.
@@
Evaluation Criteria: We use the Region-Based Detection Criterion (RBDC) and the Track-Based Detection Criterion (TBDC) as proposed in (Ramachandra and Jones, 2020) as our primary evaluation criteria and report the area under the curve (AUC) for false positive rates per frame from 0 to 1. We also report frame-level AUC (Mahadevan et al., 2010) scores. As highlighted in previous works (Ramachandra and Jones, 2020) frame-level AUC only evaluates temporal accuracy and disregards spatial localization of anomalies. RBDC and TBDC measure a method’s capacity to accurately identify anomalous spatio-temporal regions within a given video sequence.